OpenStudy (anonymous):

f(X) = x + 4, find the value of x if f(x) = 6 f(3) = 6 + 4 f(3) = 10?

OpenStudy (anonymous):

Also, If g)(x) = x^2 - 1, find g(3) 3 = x^2 - 1 4 = x^2

OpenStudy (hba):

What ? ^^

OpenStudy (anonymous):

-_-

OpenStudy (hba):

I Will Show You Wait

OpenStudy (hba):

\[g(x)=x ^{2}-1\]

OpenStudy (hba):

\[g(3)=(3)^{2}-1\]

OpenStudy (hba):

\[g(3)=9-1\]

OpenStudy (hba):

\[g(3)\]=8

OpenStudy (anonymous):

Ok, I see. Was I right on the first one?

OpenStudy (hba):

I Dont Think So

OpenStudy (anonymous):

f(X) = x + 4, find the value of x if f(x) = 6 f(x) = 6 + 4 f(x) = 10

OpenStudy (hba):



OpenStudy (anonymous):

How do I know when to replace: g(x) and when not to replace it?

OpenStudy (anonymous):

Some problems require me to get rid of g(x) and some don't.

OpenStudy (hba):

Well Its Easy See What You Are Asked

OpenStudy (anonymous):

? :P

OpenStudy (hba):

Well Math Requires Practice :)

OpenStudy (anonymous):

oo [dp

OpenStudy (hba):

What Do You Want ?

OpenStudy (anonymous):

Thank you

OpenStudy (hba):

You Are Welcome :)